OpenStudy (anonymous):

11^(6n-11)= 1/14614 how do you solve for n??

hartnn (hartnn):

its 14641 not 14614 and 14641 = 11 ^4 so u get 1/11^4 = 11^(6n-11) so 6n-11 =- 4 solve this

OpenStudy (anonymous):

lol i mix up numbers....

OpenStudy (anonymous):

and thank you muchos!

hartnn (hartnn):

welcome ^_^